### v1

@@ 5.3.1 Service Usecase
  • Quantum search is identified in a total of 2/41 studies, i.e., approximately 5% of reviewed literature [S5][S17]. For example, the study [S] develops an experimental solution to execute quantum microservives with a usecase on searching for the best path in the context of travelling salesman problem. Quantum microservice implementation is integrated into a hybrid microservices architecture, tested on Amazon Braket across three quantum hardware vendors.
  • Algorithmic execution is identified in a total of 2/41 studies, i.e., approximately 5% of reviewed literature [S6][S24]. Execution of classical or quantum specific algorithms, such as Shor’s algorithm on quantum machines is a complex problem. For example , the study [S6] proposes a solution named quantum algorithm as a service enabling developers to abstract quantum components using software as a service and function as a service, supporting algorithm deployment across multiple quantum cloud providers and hardware platforms.
@@
  • Process Automation and Optimisation is identified in a total of 4/41 studies, i.e., approximately 10% of reviewed literature [S2][S5][S7][S15]. The study [S7] proposes a hybrid quantum–classical service approach using Amazon Braket and demonstrates its applicability through a software optimization case study that tackles problems beyond the capabilities of classical algorithms.
@@
The usecase may impact the selection of programming languages and tools for service implementation. For example, the study [S3] uses Q# as the programming language that can be developed and compiled in Microsoft .Net framework for executing quantum algorithms.
5.3.2 Service Implementation
It involves programming languages that represent a system of notation or source coding scripts for implementing quantum services to manage and operationalise QC resources [25] [23]. In recent years, a number of Quantum Programming Languages (QPLs), including but not limited to Q# by Microsoft or Cirq by Google have emerged to provide specialized programming syntax and execution environment to develop, execute, and deploy quantum source code. Insights into programming languages can reveal if classical programming languages (e.g., C, Java, Python etc.) are suitable for QCaaS implementation or if there is need for more specialised QPLs (Q#, Cirq etc.). The SMS results indicate three programming languages, namely Python, Java, and Q#, as the preferred languages to implement quantum services. Based on the details of source coding, Figure 5 f) distinguishes between native code of a language and specialised libraries/application programming interfaces (APIs) being developed using a specific language. For example, the studies [S4, S5] used native Python code to implement quantum micro-servicing for experimentation. In comparison, the study [S1] used Flask as a Web framework written in Python to develop a solution for optimal delivery of quantum services on Amazon Braket. Python is the most preferred programming language both in terms of native code as well as specialised libraries of Python that include Flask and Qiskit as open-source language frameworks and Cirq which is adopted by Google. For further details on quantum programming languages, a recently conducted study [59] combines software repository mining and a practitioner survey as a mixed-method research approach to develop taxonomies of current uses and informs about emerging challenges in quantum programming as perceived by quantum developers.
  • Python as a programming language represents the predominant choice to implement quantum software services identified in a total of 19/41 studies, i.e., 46% of the total reviewed literature. Python-based service implementation can be generally classified into five sub-categories, such as native Python code [S4][S6][S35], Python Flask [S1][S5][S15], Python with Qiskit [S23][S29][S31], and Python with other languages and/or frameworks [S16][S20][S24]. For example, a quantum algorithm implemented in Python, can utilise the API Gateway pattern, for its deployment as quantum algorithm as a service on Amazon Braket platform [S6]. Python Flask is a lightweight web framework that can be used to develop web applications using Python. Flask as a web development framework for is considered as simple, flexible, and easy to get started with making it a preferred option both beginners and professionals. For example, the source code for qunatum web services is generated in Python, using the Flask Web application framework [S15].
@@ 5.3.2 Service Implementation
  • Others reprents service implementation using a language and/or implementation framework not covered in any for these languages, identified in one study [S12]. The study [S12] proposes OpenAPI is a specification for defining and documenting RESTful APIs in a standard, machine-readable format. It allows developers to describe an API’s endpoints, request/response formats, authentication, and other details, typically using JSON or YAML.
@@
L Quantum service assembly phase focuses on assembling, i.e., implementing the modelled services, that is mainly supported by: Service usecase represents the practical context/use cases of quantum services that range from quantum search, to simulation, cyber security and proof-of-the-concept implementation of quantum algorithms. Usecass also impact the selection of programming languages and tools for implementation. Service implementation involves programming languages and/or frameworks as a system of notation for source-coding of quantum services. Classical programming languages, such as Python, represent a predominant choice over recently emerged quantum programming languages (Q#) due to more comprehensive documentation and familiarity with Python in the community of service developers.
@@ 5.4 Service Deployment
As the last phase of quantum service development lifecycle, this phase supports the selection of QC platforms on which the assembled services can be deployed for their operationalisation and execution [17]. Platform providers also referred to as quantum vendors offer computing infrastructure in terms of hardware as well as software that allows service developers to develop and/or utilise the quantum services. Deployment represents the last activity in the service life cycle that can be represented with modeling notations such as the UML deployment diagram, as shown in Figure 6. This SMS identified a total of three quantum vendors for the deployment of quantum services, namely Amazon Bracket, IBM Quantum, and Rigetti. Amazon Braket (a managed Amazon Web Services (AWS)) is the most preferred platform to design, test, and run quantum algorithms. One of the reasons for selecting Amazon Braket for service deployment is that it can allow service users/developers to design their own quantum algorithms. This can be particularly useful for novice developers unfamiliar with the technicalities of quantum systems to utilise a set of pre-built algorithms, tools, and documents to develop and manage quantum services on Amazon platform.
@@
  • D-Wave Leap developed by D-Wave systems is a quantum annealing computer designed to solve complex optimization problems. It uses quantum annealing rather than gate-based computation, making it suitable for tasks like logistics, machine learning, and scheduling on specialized quantum hardware. It has been identified in a total of 3 studies [S11][S23][S41]. The study [S41] classical-quantum cloud architecture blueprint that spans from code-level development to production-ready quantum software applications, implemented on the D-Wave quantum platform for solving real-world optimization problems.
@@
L Quantum service deployment is enabled by deploying the implemented services on a QC server - enabling quantum service-orientation. Deployment involves selecting quantum computing (QC) platforms for executing services, marking the final phase in the service lifecycle. This study identified Amazon Braket, IBM Quantum, and Rigetti as major platforms for qunatum service deployment and delivery. Amazon Braket is the most preferred platform due to its accessible tools, support for custom algorithms, and ease of use for novice developers.
@@ 6 A Reference Architecture for QCaaS
In software engineering context, a reference architecture acts as a system-wide blueprint, defining computational elements as architectural components and representing their interactions through architectural connectors [60] [61]. The component and connector level architectural view can enable software designers to apply reuse knowledge and best practices and facilitates communication between domain professionals (a.k.a. system stakeholders) to implement the architecture [36]. Based on the results of RQ-1 detailed in Section 5 with summarised findings in Table 2, we applied the layered architecture pattern [62] to organise the service lifecycle activities into three layers namely service development, service deployment, and service split layers, illustrated in Figure 7 and detailed below. Figure 7 a) shows an abstract view to conceptualise the structural composition of the architecture, whereas Figure 7 b) uses the abstract view to present a fine-grained and concrete view of the architecture in terms of i) architectural layers, ii) phases and activities encapsulated inside each layer, iii) human roles, and iv) service artifacts. A fine-grained representation of the reference architecture, structured in individual layers and their encapsulated elements, is detailed below, as illustrated in Figure 7.
Tool support: To support architecture-based development, we have developed a prototype named ‘QADL: Quantum Architecture Description Language’ that provides a specification language, design space, and execution environment to architect quantum software services. QADL represents a proof-of-the-concept system that enables designers to specify architectural components (representing computational services), define component interconnections (enabling service communication) to support architecture-based development. Further details about QADL are provided in [53]
@@ 6.1 Layers of the Architecture
Architectural layering allows a separation of functional concerns which means architectural elements that support similar functionality can be structured in the same layer. For example, Figure 7 shows that conception, modelling, and assembly of quantum services can be unified into a layer named service development layer, service execution/hosting is presented as part of deployment layer, and the split between quantum and classic execution is presented as part of the service split layer. In addition to supporting the structuring of the system, a reference architecture provides a template that helps to design a (software) solution for a particular domain [60] - quantum computing in this case. Reference architectures also provides a common architecting vocabulary, such as functional requirements and their representation as architectural components and connectors. For example, the service development layer has one of the phases named Quantum Service Conception that encapsulates two activities namely Functional Specifications and Quality Attributes. These two activities allow a human role in the architecture such as Service Developer who can specify the required functionality and desired quality of the service. The phase produces a Quantum Significant Requirements (QSRs) as an artifact, which provides the foundation for service design. In short, architectural structuring as layers contains phases that encapsulate multiple activities, enabling the human role, i.e., the service developer to produce the service artifact for an incremental conception, modelling, assembly, and deployment of QCaaS.
@@ 6.2 Layering of Service Lifecycle Activities
The three architectural layers comprises of phases and activities of service lifecycle that are organised based on IBM service-oriented architecture (SOA) [26] to develop the reference architecture. For example, in Table 2, the phase referred to as modeling helps to focus on how to model or represent the quantum services for their implementation, while the activities named (i) modeling notation show class and component diagrams as UML-based modelling notation to enable service modelling and (ii) patterns indicate best practices to model the QC services. To architect QCaaS, we divided the Model activity from SOA life cycle into two activities, namely Conception and Modelling to distinguish between functional needs (conception) and representation (modelling) of quantum service design. The distinction allows a fine-grained representation of the reference architecture that delineates (i) the conception of functional needs and (ii) the model that architecturally represents the functional needs for their implementation. Model represents the conception as the design specification of functional needs for quantum services. For example, the initial row of Table 2 highlights that the conception, i.e., required functionality that enables the delivery of quantum software is modeled, i.e., architecturally represented using a UML deployment diagram and applying the API Gateway pattern. The reference architecture in Figure 7 do not have the Manage phase from SOA life cycle as we could not find any evidence in the literature that supports identity, compliance, and business metrics management of quantum services.
6.3 Human Roles in the Architecting QCaaS
@@
Human roles represent human knowledge, expertise, or involvement as part of the architecting that enables development or utilisation of the quantum services [22]. There are two types of human roles in the reference architecture, referred to as service developers and service users. Service users can be individual(s) or a group of users (part of a team or an organisation) that requires quantum computation. In contrast, service developers can include a multitude of roles that include but are not limited to quantum service developers, quantum algorithm designers, and quantum domain engineers. A recent study on architecting quantum software highlights the need for quantum-specific expertise, such as quantum software architects who can map the operations of QuBits to architectural components, and quantum code managers who can simulate and analyse the flow of quantum information processing [19]. A specific human role, such as quantum domain engineer, can analyse quantum-specific attributes like mapping between QuGates and their corresponding QuBit representation. Quantum domain engineer focuses on design and fabrication activities to improve QPU performance by optimising quantum algorithms and programming languages. Quantum domain engineers, guide quantum hardware and software teams to realise quantum significant requirements that need to be implemented as quantum algorithms for execution on quantum computing platforms, as in Figure 7 b).
@@ 6.4 Service Artifacts
Service artifacts refer to specific outcomes, such as a document, design model, or source code script that enable the development and delivery of the quantum service. As in Figure 7, each phase in the architecture layer produces an artifact as an outcome of the specific phase. There are four artifacts referred to as Quantum Significant Requirements, Quantum Service Design, Quantum Service Implementation and Quantum Service Deployment. For example, the quantum service conception phase produces QSR having function and quality attributes as an artifact that supports modelling and pattern-based design of quantum service.
Figure 7: A Layered View of the Reference Architecture for QCaaS.
6.5 Example Application Scenario for the Reference Architecture
@@
  • Phase I - Service Conception: The listings below highlight the functional specification of the quantum services to generate the prime factors of a given integer. To provide this functionality, the computations need to be split between a classical machine and a quantum machine. There is also a need to assess the correctness of the implementation and efficiency of the solution for prime factorisation. The listings below as the functional aspects and quality attributes two quality attributes concerning an efficient utilisation of the QC resources in terms of the utilisation of the QuBits. The quality attributes complement the functional specification with the desired quality of the solution.
@@
    Develop a solution that acquires an input integer N and outputs its prime factors F. - Split classical and quantum computations. - Validate correctness and efficiency.
@@
    QuBit utilisation The solution should efficiently utilise the available qubits by minimising the number of qubits required for factoring integers of a given size.
@@
    For example, Figure 8 indicates two architecting activities as part of service conception to specify the functional and quality aspects of prime factorisation. Functional and quality aspects are specified as quantum significant requirements that represents an architectural artifact of the conception phase. The service architect/developers can rely on the QSRs to utilise available modelling notations (e.g., UML, ADL) and apply any patterns (e.g., service orchestrator) to create a service design that acts as an artifact for assembling the microservices.
@@
  • Phase II - Service Modeling: The QSRs, as part of the conception, act as the foundation to create a service design model that is reflective of the functional aspects. The quantum service modeling phase relies on the following two activities.

### v2

@@ 5.3.1 Service Usecase
  • Quantum search is identified in a total of 2/41 studies, i.e., approximately 5% of reviewed literature [S5][S17]. For example, the study [S5] develops an experimental solution to execute quantum microservives with a usecase on searching for the best path in the context of travelling salesman problem. Quantum microservice implementation is integrated into a hybrid microservices architecture, tested on Amazon Braket across three quantum hardware vendors.
  • Algorithmic execution is identified in a total of 2/41 studies, i.e., approximately 5% of reviewed literature [S6][S24]. Execution of classical or quantum specific algorithms, such as Shor’s algorithm on quantum machines is a complex problem. For example , the study [S6] proposes a solution named quantum algorithm as a service enabling developers to abstract quantum components using software as a service and function as a service, supporting algorithm deployment across multiple QC providers and hardware platforms.
@@
  • Process Automation and Optimisation is identified in a total of 4/41 studies, i.e., approximately 10% of reviewed literature [S2][S5][S7][S15]. The study [S7] proposes a hybrid quantum–classical service approach using Amazon Braket and demonstrates its applicability through a software optimisation case study that tackles problems beyond the capabilities of classical algorithms.
@@
The type of use case can influence the selection of implementation frameworks, programming languages and tools for service implementation. For instance, study [S3] employs Q#, within the Microsoft .NET framework to implement and execute quantum algorithms.
5.3.2 Service Implementation Frameworks
Quantum service implementation relies on frameworks and programming languages etc., which provide the notation and source code scripts necessary to develop, manage, and operationalise QC resources [27] [25]. In recent years, several Quantum Programming Languages (QPLs) such as Q# (Microsoft) and Cirq (Google) have emerged, offering specialised syntax and execution environments for developing, executing, and deploying quantum source code. Analysing the programming languages helps determine whether classical languages (e.g., Python, Java, C) suffice for QCaaS implementation or if specialised QPLs are needed. The SMS results identify Python, Java, and Q# as the primary languages for implementing quantum services. Figure 5 f) differentiates between native code and specialised libraries or APIs built using a specific language. For example, studies [S4][S5] used native Python code for quantum microservices experimentation, while study [S1] employed Flask, a Python-based web framework, to optimise quantum service delivery on Amazon Braket. Python remains the most widely used language, both in its native form and through frameworks and libraries such as Flask, Qiskit, and Google’s Cirq. For further details on quantum programming languages, a recently conducted study [60] combines software repository mining and a practitioner survey as a mixed-method research approach to develop taxonomies of current uses and informs about emerging challenges in quantum programming as perceived by quantum source code developers.
  • Python as a programming language represents the predominant choice to implement quantum software services identified in a total of 19/41 studies, i.e., 46% of the total reviewed literature. Python-based service implementation can be generally classified into five sub-categories, such as native Python code [S4][S6][S35], Python Flask [S1][S5][S15], Python with Qiskit [S23][S29][S31], and Python with other languages and/or frameworks [S16][S20][S24]. For example, a quantum algorithm implemented in Python, can utilise the API Gateway pattern, for its deployment as quantum algorithm as a service on Amazon Braket platform [S6]. Python Flask is a lightweight web framework that can be used to develop web applications using Python. Flask as a web development framework for is considered as simple, flexible, and easy to get started with making it a preferred option both beginners and professionals. For example, the source code for qunatum web services is generated in Python, using the Flask Web framework [S15].
@@ 5.3.2 Service Implementation Frameworks
  • Others reprents service implementation using a language and/or implementation framework not covered in any of the above-mentioned languages, identified in one study [S12]. The study [S12] proposes OpenAPI is a specification for defining and documenting RESTful APIs in a standard, machine-readable format. It allows developers to describe an API’s endpoints, request/response formats, authentication, and other details, typically using JSON or YAML.
@@
L Quantum service assembly phase focuses on assembling, i.e., implementing the modelled services, that is mainly supported by: Service usecase representing the practical context or a usage scenario of quantum services that range from quantum search, to simulation, cyber security and proof-of-the-concept implementation of quantum algorithms. Usecass also impact the selection of programming languages and tools for implementation. Service implementation involves programming languages and/or frameworks as a system of notation for source-coding of quantum services. Classical programming languages, such as Python, represent a predominant choice over recently emerged quantum programming languages (Q#) due to more comprehensive documentation and familiarity with Python in the community of service developers.
@@ 5.4 Service Deployment
As the final phase of the quantum service development lifecycle, the deployment phase focuses on selecting QC platforms for operationalising and executing assembled services [17]. Platform providers, or quantum vendors, supply both hardware and software infrastructure that enables service developers to build and utilise quantum services. Deployment can be represented using modeling notations such as UML deployment diagrams, as illustrated in Figure 6. The SMS identified three primary quantum vendors for service deployment: Amazon Braket, IBM Quantum, and Rigetti. Among these, Amazon Brake, a managed service on AWS, is the most preferred platform for designing, testing, and running quantum algorithms. Its popularity stems from its flexibility, allowing developers, including novices, to design their own quantum algorithms or leverage pre-built algorithms, tools, and documentation to develop and manage quantum services on the platform.
@@
  • D-Wave Leap developed by D-Wave systems is a quantum annealing computer designed to solve complex optimization problems. It uses quantum annealing rather than gate-based computation, making it suitable for tasks like logistics, machine learning, and scheduling on specialized quantum hardware. It has been identified in a total of 3 studies [S11][S23][S41]. The study [S41] presents a quantum-classic based cloud architecture as a blueprint that spans from code-level development to production-ready quantum software, implemented on the D-Wave quantum platform for solving real-world optimisation problems.
@@
L Quantum service deployment is enabled by deploying the implemented services on a QC server - enabling quantum service-orientation. Deployment involves selecting quantum QC platforms for executing services, marking the final phase in the service lifecycle. This study identified Amazon Braket, IBM Quantum, and Rigetti as major platforms for quantum service deployment and delivery. Amazon Braket is the most preferred platform due to its accessible tools, support for custom algorithms, and ease of use for novice developers.
@@ 6 A Reference Architecture for QCaaS
In software engineering context, a reference architecture acts as a system-wide blueprint, defining computational elements as architectural components and representing their interactions through architectural connectors [61] [62]. The component and connector level architectural view can enable software designers to apply reuse knowledge and best practices and facilitates communication between domain professionals (a.k.a. system stakeholders) to implement the architecture [38]. Based on the findings of RQ-1 presented in Section 5 and summarised in Table 2, we applied the layered architecture pattern [63] to organise quantum service lifecycle activities into three layers: service development, service deployment, and service split, as illustrated in Figure 7. Figure 7 a) provides an abstract view to conceptualize the structural composition of the architecture, while Figure 7 b) translates this abstract view into a more detailed, concrete representation of the architecture in terms of i) architectural layers, ii) phases and activities encapsulated inside each layer, iii) human roles with tool support, and iv) service artifacts. A fine-grained representation of the reference architecture, structured in individual layers and their encapsulated elements, is detailed below and visually represented in Figure 7.
Tool support for the reference architecture: To support architecture-based development, we have developed a prototype named ‘QADL: Quantum Architecture Description Language’ that provides a specification language, design space, and execution environment to architect quantum software services. QADL represents a proof-of-the-concept system that enables designers to specify architectural components (representing computational services), define component interconnections (enabling service communication) to support architecture-based development. Further details about QADL are provided in [54]
@@ 6.1 Layers of the Architecture
Architectural layering enables the separation of functional concerns, allowing elements that support similar functionality to be grouped within the same layer. For example, in Figure 7, the service development layer unifies the conception, modeling, and assembly of quantum services; the deployment layer covers service execution and hosting, and the service split layer handles the division between quantum and classical execution. Beyond structuring the system, a reference architecture provides a template for designing software solutions within a specific domain, in this case, quantum computing in general with a specific focus on quantum service-orientation [61]. It also establishes a common architectural vocabulary, linking functional requirements to architectural components and connectors. For example, within the service development layer, the Quantum Service Conception phase includes two key activities: Functional Specifications and Quality Attributes. These activities involve human roles, such as the service developer, who defines the required functionality and desired quality of the service. The outcome of this phase is a Quantum Significant Requirements (QSRs) artifact, which forms the foundation for service design. In summary, layered architectural structuring organises phases containing multiple activities, enabling human roles to incrementally produce service artifacts through conception, modeling, assembly, and deployment of QCaaS.
@@ 6.2 Layering of Service Lifecycle Activities
The three architectural layers encompass the phases and activities of the service lifecycle, organized according to IBM’s service-oriented architecture (SOA) [28] to guide the development of the reference architecture. For example, in Table 2, the Modeling phase focuses on how to represent quantum services for implementation. Within this phase, the activities include: (i) modeling notation, which uses UML-based class and component diagrams to enable service modeling, and (ii) patterns, which provide best practices for modeling QC services. To better structure QCaaS, we split the ‘Model’ activity from the SOA lifecycle into two distinct activities: ‘Conception’ and ‘Modeling’. This separation distinguishes functional needs (conception) from their architectural representation (modeling). Conception defines the design specifications for the functional requirements of quantum services, while Modeling represents these requirements architecturally for implementation. For example, the first row in Table 2 shows that the required functionality enabling quantum software delivery is architecturally represented using a UML deployment diagram and the API Gateway pattern. It is vital to mention that the reference architecture in Figure 7 omits the Manage phase from the SOA lifecycle, as the literature provides no evidence supporting identity, compliance, or business-metrics management for quantum services. The SMS process could not find any evidence in the literature that supports identity, compliance, and business metrics management of quantum services.
6.3 Human Roles and Tools to Architect QCaaS
@@
Human roles in the reference architecture capture the knowledge, expertise, and involvement required to develop or utilise quantum services [24]. Two primary types of roles are identified: service developers and service users. Service users may be individuals or teams within an organisation that require quantum computation. In contrast, service developers encompass various specialised roles, including quantum service developers, quantum algorithm designers, and quantum domain engineers. Recent studies emphasise the importance of quantum-specific expertise, such as quantum software architects, who can map QuBit operations to architectural components, and quantum code managers, who simulate and analyze quantum information flow [19]. A quantum domain engineer, for example, focuses on analysing quantum-specific attributes, such as the mapping between QuGates and their corresponding QuBit representations. Quantum domain engineers can provide input to the design and fabrication of quantum processing units (QPUs), optimising algorithms and programming languages to enhance performance. Quantum domain engineers also guide hardware and software teams in realizing Quantum Significant Requirements (QSRs), implementing them as quantum algorithms for execution on quantum computing platforms, as illustrated in Figure 7 b).
In addition to human roles, tool support plays a crucial role in enabling automation and user-driven customisation of architecting QCaaS [64]. Tools such as Qiskit, Cirq facilitate circuit design, compilation, and optimisation, while platforms like IBM Quantum Composer, Amazon Braket, and Azure Quantum enable service orchestration and hybrid quantum–classical integration. These tools can enable automation of tasks such as QuBit utilisation, circuit design, resource allocation, and workflow execution to reduce manual intervention and improve efficiency of the architecting process [65]. Tools can also provide customisation and adaptability through inputs from human roles such as quantum service developers, who define algorithms, and system architects, who specify service configurations and quality requirements. This synergy between tool-driven automation and human-provided input ensures that QCaaS development can be efficient with humanăoversight and supervision.
@@ 6.4 Service Artifacts
Service artifacts are tangible outcomes such as documents, design models, or source code, that support the development and delivery of quantum services. In Figure 7, each phase within the architectural layers produces a specific artifact. Four key artifacts are identified: QSRs, Quantum Service Design, Quantum Service Implementation, and Quantum Service Deployment. For instance, the Quantum Service Conception phase generates QSRs, capturing functional and quality attributes that guide the modeling and pattern-based design of the quantum service.
Figure 7: Reference Architecture for QCaaS - a Layered View.
6.5 Example Scenario for the Reference Architecture
@@
  • Phase I - Service Conception: The listings below outline the functional specifications for quantum services designed to generate the prime factors of a given integer. To achieve this functionality, computations are divided between a classical compute and a quantum compute. Additionally, it is necessary to evaluate both the correctness of the implementation and the efficiency of the prime factorisation solution. The listings capture the functional aspects alongside two quality attributes, focusing on the efficient utilisation of QC resources, particularly the effective use of QuBits. Quality attributes complement the functional specifications by defining the desired characteristics and performance standards of the solution.
@@
    Design solution for inputting the integer N and outputting prime factors F. - Split the solution for classical and quantum hybrid computations. - Validate correctness of solution and efficiency of computations.
@@
    QuBit utilisation The designed solution to efficiently utilise QuBits used for factoring integers of the given size.
@@
    For example, Figure 8 illustrates two architecting activities within the service conception phase, focusing on specifying the functional and quality aspects of prime factorisation. These aspects are captured as Quantum Significant Requirements (QSRs), serving as an architectural artifact of the conception phase. Service architects and developers can use the QSRs to apply appropriate modeling notations (e.g., UML, ADL) and patterns (e.g., service orchestrator) to create a service design, which acts as an artifact for assembling the microservices.
@@
  • Phase II - Service Modeling: The QSRs, produced during the conception phase, serve as the foundation for creating a service design model that reflects the functional aspects of the quantum service. The quantum service modeling phase builds on this foundation through the following two activities.
    A. Modelling notation for quantum services: